# Is Quadratic Gravity ghost free because the ghost is only virtual?

### v1

@@ 2.1 The free Stelle action in the De Donder gauge
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h-4\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$2\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h+2\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h-\kappa^{2}(\beta-\frac{\alpha}{2})\square^{2}h=0.$
Simplifying terms
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{4}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h-\kappa^{2}(\beta-\frac{\alpha}{2})\square^{2}h=0.$ (2.14)
Therefore only $h^{1}$ may survive, as for $h^{2}$ and $h^{3}$ the last gives an algebraic equation for these modes with solution zero. For instance, for the mode corresponding to $m^{2}_{2}$ one finds
$\frac{2}{3}\bigg(1-\frac{2(3\beta-\alpha)}{\alpha}\bigg)m_{2}^{2}h^{2}+\kappa^{2}(\beta-\frac{\alpha}{2})m_{2}^{2}h^{2}=0.$
The solution is $h^{2}=0$. The same conclusion goes for $h_{3}$, which is $h_{3}=0$. The massive modes are then given by (2.9). For the massless mode, one still has to satisfy the De Donder gauge (2.4). But this situation is known in GR that it constraints the value $h^{1}$ to vanish. To see this, note that the De Donder condition for the polarizations lead to
$2k^{\mu}\epsilon^{1}_{\mu\nu}=k_{\nu}\epsilon^{1}$
This eliminates $4$ polarizations, leading to $6$ independent ones. A further gauge transformation
$h_{\mu\nu}\to h_{\mu\nu}+\partial_{\mu}\xi_{\nu}+\partial_{\nu}\xi_{\mu}.$
This eliminates $4$ adeitional degrees of freedom. The vector $\xi_{\nu}$ can be chosen such that the polarization is (2.8), see for instance formulas (2.37) to (2.43) of reference [158]. Therefore $h_{1}$ can be chosen to be zero. But then $h=0$ and the equation (2.13) reduces to
$\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}=0.$
Therefore the equation (2.11) has been found again, implying that $h^{3}_{\mu\nu}=0$.
One may include a more general argument. The equation (2.14) is an equation for the trace of the form
$(1-\frac{\square}{m_{4}^{2}})\square h=0,$
where $m^{2}_{4}$ is a new scale. Then take $\partial^{\mu}$ to the equation (2.13), employ the De Donder gauge in the first term, go to the momentum space and multiply by $k^{\nu}$. Then a relation analogous to (2.7) will be found namely
$F(m_{i}^{2})k_{\mu}\epsilon_{4}=0,$ (2.15)
with $F(m_{i}^{2})$ a function of all the mass scales. Then $\epsilon_{4}=0$ and only the massless mode survives, which the De Donder gauge converts its polarization into a traceless and ”orthogonal” to the propagation direction.
The above conclusion shows that the third mode can not be turned on as a ”plane wave”. However, this does not forbid this term to generate Yukawa like potentials $\frac{e^{-m_{3}r}}{r}$ in spherically symmetric solutions of gravity [32]-[35]. In some sense, the absence of this mode in the free modes imitates confinement. On the other hand, the mode can be seen in classical spherical solutions. It is not clear yet for the author to refer this mode as confined or virtual, or even to employ a new terminology.
One may argue that, at quantum level, the third mode should not be eliminated. Instead, one should employ the Gupta-Bleuler quantization, which imposes weaker conditions at operator level. It will be shown below that nevertheless, even taking into account this quantization, the third mode disappears. This observation raises a very subtle detail. Consider for instance the formula (4.74) for the quartic Pais-Uhlenbeck scalar field $\phi$ given below. These commutator of the quantum field $\phi$ defines the propagator. Here it does not. This is simple to see. If the graviton is quantized and the third mode is absent (as it will be shown below at quantum level), then $\hat{h}_{\alpha\beta}$ does not contain $m_{3}^{2}$. Neither does the commutator
$\hat{h}_{\mu\nu},\hat{h}_{\alpha\beta}]=F(m_{2}^{2},k)$
However, the wave operator and therefore the gauge fixed lagrangian leading to the equations (2.13) characterizing the graviton does depend on the mass scale $m_{3}^{2}$. Therefore
$\hat{h}_{\mu\nu},\hat{h}_{\alpha\beta}]\neq D_{\alpha\beta,\mu\nu},$ (2.16)
where $D_{\alpha\beta,\mu\nu}$ is the propagator of the model. Situations such as (4.74) are lucky, not the general rule. This observation will be fundamental in the analysis below.
The next subsection calculates the propagator on the right hand side of (2.16), thereby confirming that there is no equality between the left and right hand side of this formula.
@@ 2.2 The free action in the Stelle gauge
$P^{2}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\theta_{\alpha\nu}+\theta_{\beta\nu}\theta_{\alpha\mu})-\frac{1}{3}\theta_{\beta\alpha}\theta_{\mu\nu}$
$P^{1}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\omega_{\alpha\nu}+\theta_{\beta\nu}\omega_{\alpha\mu}+\theta_{\alpha\mu}\omega_{\beta\nu}+\theta_{\alpha\nu}\omega_{\beta\mu})$
@@
$\omega_{\mu\nu}=\frac{\partial_{\mu}\partial_{\nu}}{\square},\qquad\theta_{\mu\nu}=\eta_{\mu\nu}-\frac{\partial_{\mu}\partial_{\nu}}{\square}.$
@@
$S=\frac{2}{\kappa^{2}}\int h^{\mu\nu}\square\bigg[\frac{1}{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)P^{2}-\bigg(1+\frac{\kappa^{2}(\alpha-3\beta)}{2}\square\bigg)P^{0-s}$
$\bigg(\frac{1}{2}+\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}\bigg]_{\alpha\beta,\mu\nu}h^{\alpha\beta}d^{4}x.$ (2.22)
@@
$S=\frac{2}{\kappa^{2}}\int h^{\mu\nu}\square\bigg[\frac{1}{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)P^{2}-\bigg(1+\frac{\kappa^{2}(\alpha-3\beta)}{2}\square\bigg)P^{0-s}$
$\bigg(\frac{1}{2}+\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}+\frac{\kappa^{4}}{8\lambda}\square[P^{1}-P^{0-w}]\bigg]_{\alpha\beta,\mu\nu}h^{\alpha\beta}d^{4}x.$

### v4

@@ 2.1.1 The Gupta-Bleuler quantization scheme
$\sum_{\lambda}\epsilon_{2}^{\lambda}b_{\lambda}(k)|\Psi>=0.$ (2.12)
The absence of the mode with $m_{2}^{2}$ in the average leads to
The independence between the modes $1$ and $3$ leads to
For the third mode one has from (2.10) that
$\sum_{\lambda}\epsilon^{3\lambda}_{\mu\nu}c_{\lambda}(k)||\Psi>=\sum_{\lambda}\frac{1}{3}\bigg(\eta_{\mu\nu}+\frac{k_{\mu}k_{\nu}}{m_{3}^{2}}\bigg)\epsilon^{\lambda}c_{\lambda}|\Psi>$
Quantization should be performed by considering general polarizaion and assuming that the physical states are constrained by the above conditions.
@@ 2.2 The free action in the Stelle gauge
$P^{2}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\theta_{\alpha\nu}+\theta_{\beta\nu}\theta_{\alpha\mu})-\frac{1}{3}\theta_{\beta\alpha}\theta_{\mu\nu},$
$P^{1}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\omega_{\alpha\nu}+\theta_{\beta\nu}\omega_{\alpha\mu}+\theta_{\alpha\mu}\omega_{\beta\nu}+\theta_{\alpha\nu}\omega_{\beta\mu}),$
@@
$\omega_{\mu\nu}=\frac{\partial_{\mu}\partial_{\nu}}{\square},\qquad\theta_{\mu\nu}=\eta_{\mu\nu}-\frac{\partial_{\mu}\partial_{\nu}}{\square},$
@@
$S=\frac{2}{\kappa^{2}}\int h^{\mu\nu}\square\bigg[\frac{1}{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)P^{2}-\bigg(1-\frac{\kappa^{2}(\alpha-3\beta)}{2}\square\bigg)P^{0-s}$
$\bigg(\frac{1}{2}-\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}\bigg]_{\alpha\beta,\mu\nu}h^{\alpha\beta}d^{4}x.$ (2.19)
@@
$S=\frac{2}{\kappa^{2}}\int h^{\mu\nu}\square\bigg[\frac{1}{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)P^{2}-\bigg(1-\frac{\kappa^{2}(\alpha-3\beta)}{2}\square\bigg)P^{0-s}$
$\bigg(\frac{1}{2}-\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}+\frac{\kappa^{4}}{8\lambda}\square[P^{1}-P^{0-w}]\bigg]_{\alpha\beta,\mu\nu}h^{\alpha\beta}d^{4}x.$